OpenStudy (faman39):

Kenny goes shopping for school clothes during a tax−free weekend, and he saves 8% sales tax. Use a percent proportion to determine how much he would save if he bought clothes worth $200 . Which of the following is true?

OpenStudy (faman39):

wait lemme give u the choices of worksheet

OpenStudy (wasiqss):

ook

OpenStudy (faman39):

The solution is $16 , and is a denominator in the proportion. The solution is $16, and is a numerator in the proportion. The solution is 2,500, and is a denominator in the proportion. The solution is 2,500, and is a numerator in the proportion.

OpenStudy (wasiqss):

answer is b

OpenStudy (experimentx):

if he paid 200$ then x-8%of x = 200 or x=200/0.92 = 217.4 so the he saves 17.4$$$ But it seems the cloth was worth $200 so he did not have to pay $200 so he saves 8% of 200 = 16$

Directrix (directrix):

The solution is $16, and is a numerator in the proportion ---------------------------------------------- The question asks for a percent proportion solution technique: 8/100 = x/200 where x is the amount saved x = $16